# Semi-Supervised Object Detection: A Survey on Progress from CNN to Transformer

### v1

@@ 4 Semi Supervised Strategies
4.1 STAC
4.2 Humble Teacher
4.3 Instant-Teaching
4.4 Soft Teacher
4.5 Unbiased Teacher
4.6 ACRST
4.7 Combating Noise
4.8 MUM
4.9 ISTM
A Semi-Supervised Object Detection technique known as Interactive Self-Training with Mean Teachers (ISTM) [161] introduces an approach to rectify the oversight of inconsistencies among detection outcomes in the same image across various training iterations, as shown in Fig. 11. By utilizing non maximum suppression [106] to combine detection outcomes from different iterations and employing multiple detection heads to offer complementary information, this approach boosts the stability and quality of pseudo labels. Moreover, the incorporation of the mean teacher model [133] prevents overfitting [150] and aids in the transfer of knowledge between detection heads.
4.10 CrossRectify
4.11 SED
4.12 Label Match
4.13 DTG-SSOD
4.14 Rethinking Pse
4.15 CSD
4.16 PseCo
4.17 Active Teacher
4.18 SCMT
4.19 Semi-DETR
4.20 Sparse Semi-DETR
4.21 Omni-DETR
@@ 4.27 Consistent-Teacher
Inconsistent pseudo labels [91, 92, 93] in Semi-Supervised Object Detection (SSOD) pose a challenge that Consistent-Teacher [22] addresses. These pseudo labels introduce noise into the student’s training process, which causes serious overfitting [150] problems and compromises the construction of accurate detectors. As represented in Fig. 29, Consistent-Teacher introduces a 3D feature alignment module (FAM- 3D) [131], the Gaussian Mixture Model (GMM), and adaptive anchor assignment (ASA) [129, 130] as a strategy to minimize this issue. These components enhance the quality of the pseudo-boxes, dynamically modify the threshold values, and stabilize the pseudo-box matching with anchors.
@@ 6.2 Comparison
| SED [16] | Improves Semi-Supervised Object Detection by enforcing scale consistency and self-distillation. | Reliance on regularization and self-distillation may pose challenges for parameter tuning and scalability. |
@@
| Rethinking Pse [19] | Certainty-aware pseudo labels improve performance by addressing localization precision and class imbalance issues | Implementing certainty-aware pseudo labeling may introduce additional computational complexity during training. |
| CSD [5] | Leverages consistency constraints for both classification and localization, enhancing object detection performance using unlabeled data. | Performance of CSD may be sensitive to the effectiveness of background elimination method, impacting its generalizability. |
| PseCo [] | Enhances SSOD by integrating object detection attributes into pseudo labeling along with consistency training, leading to superior performance and faster convergence. | Techniques may add complexity and computational overhead, posing challenges in resource-constrained environments. |
@@
| Omi-DETR [3] | utilize diverse weak annotations to enhance performance and annotation efficiency. | effectiveness on larger datasets is uncertain, and its simplified annotation process may raise concerns about potential misuse. |
| One Teacher [27] | Improves SSOD on YOLOv5, tackling issues like low-quality pseudo-labeling. | Effectiveness might be impacted by the complexity of implementing YOLOv5 and its training techniques. |
| Unbiased Teacher v2[179] | Expands the applicability of SSOD to anchor-free detectors, improving performance across various benchmarks. | Challenges remain in scaling the method to large datasets and addressing domain shift issues. |
| S4OD[25] | Dynamically adjusts pseudo-label selection to balance quality and quantity, enhancing single-stage detector performance | Lack of empirical comparison with existing pseudo-label selection methods may limit the assessment of the proposed DSAT strategy’s superiority. |

### v2

@@ 4.6 Consistent-Teacher
Inconsistent pseudo labels [91, 92, 93] in Semi-Supervised Object Detection (SSOD) pose a challenge that Consistent-Teacher [22] addresses. These pseudo labels introduce noise into the student’s training process, which causes serious overfitting [150] problems and compromises the construction of accurate detectors.
As represented in Fig. 8, Consistent-Teacher introduces a 3D feature alignment module (FAM- 3D) [131], the Gaussian Mixture Model (GMM), and adaptive anchor assignment (ASA) [129, 130] as a strategy to minimize this issue. These components enhance the quality of the pseudo-boxes, dynamically modify the threshold values, and stabilize the pseudo-box matching with anchors.
4.7 Rethinking Pse
4.8 CSD
4.9 STAC
4.10 Humble Teacher
4.11 Instant-Teaching
4.12 ISMT
A Semi-Supervised Object Detection technique known as Interactive Self-Training with Mean Teachers (ISMT) [161] introduces an approach to rectify the oversight of inconsistencies among detection outcomes in the same image across various training iterations, as shown in Fig. 14. By utilizing non maximum suppression [106] to combine detection outcomes from different iterations and employing multiple detection heads to offer complementary information, this approach boosts the stability and quality of pseudo labels. Moreover, the incorporation of the mean teacher model [133] prevents overfitting [150] and aids in the transfer of knowledge between detection heads.
4.13 Combating Noise
4.14 Soft Teacher
4.15 Unbiased Teacher
4.16 DTG-SSOD
4.17 MUM
4.18 Active Teacher
4.19 PseCo
4.20 CrossRectify
4.21 Label Match
4.22 ACRST
4.23 SED
4.24 SCMT
4.25 Omni-DETR
4.26 Semi-DETR
4.27 Sparse Semi-DETR
@@ 6.2 Comparison
| SED [16] | Improves Semi-Supervised Object Detection by enforcing scale consistency and self-distillation. | Reliance on the IoU threshold criterion, which could not be optimal for all detectors and situations, and its limited benefits from multi-scale testing |
@@
| Rethinking Pse [19] | Certainty-aware pseudo labels improve performance by addressing localization precision and class imbalance issues | Implementing certainty-aware pseudo labeling add additional computational complexity during training. |
| CSD [5] | Leverages consistency constraints for both classification and localization, enhancing object detection performance using unlabeled data. | It shows less performance improvement in two-stage detectors compared to single-stage detectors. |
| PseCo [4] | Enhances SSOD by integrating object detection attributes into pseudo labeling along with consistency training, leading to superior performance and faster convergence. | Its potential struggle with generalization across diverse datasets due to variability in pseudo-label quality. |
@@
| One Teacher [27] | Improves SSOD on YOLOv5, tackling issues like low-quality pseudo-labeling. | Lowering the threshold for pseudo-labeling due to noisy pseudo-labeling in one-stage detection makes it difficult to maximize the effectiveness of one-stage teacher-student learning. |
| Unbiased Teacher v2[179] | Expands the applicability of SSOD to anchor-free detectors, improving performance across various benchmarks. | Challenges remain in scaling the method to large datasets, integrating localization uncertainty estimation for boundary prediction with the relative thresholding mechanism, and addressing domain shift issues. |
| S4OD[25] | Dynamically adjusts pseudo-label selection to balance quality and quantity, enhancing single-stage detector performance | DSAT strategy’s increased time cost is due to F1-score computation, and using the CPU version of NMS for uncertainty computation slows down training. |
| Omni-DETR [3] | utilize diverse weak annotations to enhance performance and annotation efficiency. | Effectiveness on larger datasets is uncertain, and its simplified annotation process could raise concerns about potential misuse. |